### v2

# SIM2VR: Towards Automated Biomechanical Testing in VR

Abstract.
Automated biomechanical testing has great potential for the development of VR applications, as initial insights into user behaviour can be gained in silico early in the design process. In particular, it allows prediction of user movements and ergonomic variables, such as fatigue, prior to conducting user studies. However, there is a fundamental disconnect between simulators hosting state-of-the-art biomechanical user models and simulators used to develop and run VR applications. Existing user simulators often struggle to capture the intricacies of real-world VR applications, reducing ecological validity of user predictions. In this paper, we introduce sim2vr, a system that aligns user simulation with a given VR application by establishing a continuous closed loop between the two processes. This, for the first time, enables training simulated users directly in the same VR application that real users interact with. We demonstrate that sim2vr can predict differences in user performance, ergonomics and strategies in a fast-paced, dynamic arcade game. In order to expand the scope of automated biomechanical testing beyond simple visuomotor tasks, advances in cognitive models and reward function design will be needed.
biomechanical simulation, interaction design, VR simulation alignment, automated testing, virtual reality, VR development, VR application, deep reinforcement learning
ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Systems and tools for interaction designccs: Human-centered computing Virtual realityccs: Human-centered computing User models
Figure 1. Simulating interactive user behaviour in VR typically requires re-implementing the game dynamics, logic and graphics of the VR application in another simulation engine, which is time-consuming, redundant and error-prone (left).
sim2vr (middle) enables VR designers, for the first time, to run state-of-the-art biomechanical user simulations directly in the same VR application real users are interacting with (right). By aligning simulated and real interaction with respect to VR input and output, as well as application dynamics, sim2vr allows for more ecologically valid predictions of user performance, effort and strategies, supporting developers in the early stages of the design process.
[…]
Three boxes illustrating the simulated interaction loops without SIM2VR and with SIM2VR together with the real interaction loop. The ”Without SIM2VR” loop (left) shows a biomechanical model with one eye (”Simulated User”) and three targets in the same environment (”Simulated Environment”). The ”With SIM2VR” loop (middle) shows a biomechanical model equipped with VR controller and HMD (”Simulated User”), a ”VR Application” block, and the SIM2VR logo in between. An orange frame highlights this loop. The ”Real Interaction” loop (right) shows a real user (”User”) and the same ”VR Application” block.
1. Introduction
Recent advances in deep reinforcement learning (deep RL) and biomechanical models have enabled the creation of simulated users for modelling how humans may behave in interaction tasks (Nakada18; Fischer21; Ikkala22; caggiano2022myosuite). Using advanced biomechanical models that capture the intricacies of human body movement beyond linked-segment models, such simulated users allow prediction of time-continuous movements and effort-related quantities such as muscle energy expenditure (Umberger03) and fatigue (Xia08; Rashedi15). In HCI, simulated users have been recently demonstrated for various tasks, including pointing (Klar22; Cheema20), typing (Hetzel21; Jokinen21), tracking (Fischer21), steering and choice reaction (Ikkala22). This modelling approach is promising, as it enables designers and developers to pose “what-if?” questions to assess how changes in design choices affect users’ performance and ergonomics.
This bears particular promise for HCI domains where human data is scarce or costly to collect, such as in virtual reality (VR) interactions. In VR development, testing prototypes typically involves wearing a headset, standing up and moving around in space, which incurs significant overhead and renders iterative testing slow and cumbersome (liu2023challenges). Biomechanical user simulations show significant potential to remedy this as a means of automated testing. In particular, they allow VR designers to evaluate and compare prototypes without the need for expensive user studies, which can be lengthy and stressful, especially for vulnerable participants. By using an appropriate range of different user models, a simulation-based approach can increase diversity during the testing process, thereby promoting ability-based design (murraysmith2022simulation).
To realise the potential of simulated users, we need to address the challenge of integrating model-based simulations into VR development workflows. State-of-the-art biomechanical models are not readily available in commonly used VR development environments like Unity, [^1] https://unity.com/ or they cannot be combined with deep RL approaches to generate forward simulations in those environments. For instance, while there exists a MuJoCo Unity plugin [^2] https://github.com/google-deepmind/mujoco/tree/main/unity that provides access to biomechanical models implemented in MuJoCo (Todorov12), an efficient physics engine suitable for biomechanical simulation, the models can only be controlled manually using the Unity Editor.
As a result, VR designers who wish to use automated testing must recreate the dynamics and graphics of the VR application within the simulation engine used for the biomechanical model. This, however, is impractical for three reasons. First, it introduces implementation overhead, in that two VR environments need to be maintained and updated in parallel. Second, building the replica is often tedious and cumbersome, and requires detailed knowledge of the subtleties of the considered application. Third, the resulting simulation is unlikely to create ecologically valid predictions of interactive user behaviour, as the replica will typically correspond to a low-fidelity version of the original VR application. Such an approach thus considerably complicates the reduction of differences between the simulated interaction loop and the real interaction loop, which can be described by means of several alignment problems, as detailed in Section 3. Simulations that do not adequately address these problems will provide inferior predictions of how real users interact with VR interfaces or techniques, limiting the validity of in silico testing and compromising the comparative analysis of design choices.
To overcome these limitations, we propose sim2vr, a system that integrates biomechanical user simulations directly into a given Unity application. Alignment is improved by simulating hardware within the physics engine used for biomechanical simulation, tracking virtual VR input and output using OpenXR, accessing all relevant application data directly in Unity, and synchronising the user simulation and the VR application. From a research perspective, this enables a more accurate and reliable development and validation of user simulation models by ensuring that real and simulated users interact with exactly the same virtual environment (see Figure 1). From an interactive software development perspective, our work provides a new solution to the long-standing problem of automated testing of VR interaction.
To summarise, this paper contributes
  • sim2vr, a system that for the first time integrates biomechanical simulations directly into VR application development (Section 4);
  • an empirical assessment of sim2vr’s predictive capabilities for automated testing (Section 5);
  • a critical discussion on the scope as well as current and inherent limitations of sim2vr (Section LABEL:sec:discussion-limitations).
Open-source code of sim2vr is available at https://github.com/fl0fischer/sim2vr.
2. Background and Context
Below, we situate our work in relation to previous efforts in both automated testing in VR development, and computational user modelling and simulation.
2.1. VR Development and Automated Testing
Because of the effort and time needed to don and doff equipment during prototype development, iterative testing and development in VR has a higher cost than traditional interface design for desktop or mobile software, making it typically slow and cumbersome (liu2023challenges). This underlines the need to automate testing as much as possible. Automated testing allows prototypes to be automatically evaluated and compared without the need for expensive user studies (bierbaum2003automated; karakaya2022automated), by predicting performance estimates such as completion times and error rates.
While the need for automated testing has been clear for decades, a recent study of over 300 VR projects found that 79% of the projects did not use any automatic tests, and that test automation followed common software testing practices without considering the user’s body movements, focusing on aspects such as evaluating the correctness of a response after an event was triggered (rzig2023virtual). So far, even automated testing approaches designed specifically for VR do not emulate or simulate the user’s moving body. Instead, they focus on moving and rotating the viewpoint to inspect VR scenes (wang2022vrtest), or operate on higher-level action events such as “click UI button” or “grab object” which abstract away the details of the user’s movements (bierbaum2003automated; harms2019automated; karakaya2022automated). In the domain of Natural User Interfaces (NUIs), a generative motion model was proposed for automatic testing of gestural interaction in (hunt2014automatic), but unlike our work, the model was not embedded in a learning loop; instead, it generated random movement sequences without the ability to adapt to the tested interface.
2.2. Computational User Modelling and Simulation
2.2.1. Simulation in HCI
As discussed in (murraysmith2022simulation), user-centred design could move to a more rigorous, safer and predictable process via a simulation-based approach to interaction design. The same paper highlights general potential benefits of simulation in HCI for efficiency of design, safety, supporting diversity and transparency. Simulations can be used to predict task performance in interaction, such as time taken to finish the task and how often tasks can be successfully accomplished. For example, very early work on keystroke-level models in (card1983psychology), and more recent work (kristensson2021design) used modelling to replace extensive experimentation, optimising text entry system parameters.
The general tools for simulation are improving rapidly, opening up new opportunities. As pointed out in (lavin2021simulation), we have seen recent developments in probabilistic, differentiable programming, high-performance computing and causal modelling. This, combined with the rapidly improving ability of machine learning to emulate complex aspects of human perception and behaviour (e.g. deep neural nets can simulate human-level visual perception (he2016deep), speech recognition (graves2013speech) and control (mnih2015human)), means that these new tools for simulation have an increased potential to be efficiently and usefully applied to domains such as HCI.
2.2.2. Biomechanical User Simulation
Biomechanical models and computer-based simulations are well-established (Ayoub1974). Increases in computational power have allowed them to evolve from simple models, limited to computation of mechanical loads in static postures (Winter1984), to more physiologically-accurate musculoskeletal models (Delp07; Damsgaard06). Biomechanical models were typically used for inverse simulation, e.g. using the OpenSim ecosystem (SETH2011212). Such inverse simulation methods, namely inverse kinematics, inverse dynamics and static optimisation or computed muscle control, allow the estimation of mechanical loads within the human musculoskeletal system, and neural controls of the muscles given motion tracking data of a given user’s movement as input  (Delp07). In HCI, inverse dynamics have been used to identify fundamental ergonomics issues, such as “gorilla arm” (jang2017modeling). Forward simulation methods, in contrast, are less frequently used in standalone situations, as they require muscle controls as inputs, which are extremely complex to measure experimentally. However, these have become more useful when used in conjunction with computational controllers (lee2016generating; Dembia2021; schumacher2023deprl) and efficient physics engines (Klar22), which we will exploit in this paper.
2.2.3. Optimal Feedback Control and Reinforcement Learning Models
Control-theoretic models of human motion for mouse movements were developed in (muller2017control) and supplemented by the assumption of Optimal Feedback Control (Todorov02) in (Fischer22). In such optimisation-based forward simulations, users are assumed to continuously interact with their environment by selecting controls such that a task-specific utility (e.g., a game score) is maximised, given the perceptual information available at decision time. By including intermittent control behaviour typical of humans, simulations could also capture more realistic models of human motor variability (martin2021intermittent).
Reinforcement learning has recently emerged as a suitable framework for modelling human behaviour in a flexible way: one only needs to define the states, actions and rewards, and then use RL methods to estimate the optimal policy (chen2015emergence; chen2017cognitive). In some cases, the reward function and other parameters can be inferred from human data (Banovic16; Kangasraasio17). If the reward function and state-action space, including their key limitations, are similar to those of a human, increasingly human-like behaviour has been shown to emerge through learning (oulasvirta2022computational).
Recent applications in HCI include models of typing, menu selection, multitasking and visual decision-making (oulasvirta2022computational). Using the assumptions of signal-dependent control noise and movement time minimisation, Fischer et al. (Fischer21) have shown that an RL agent can generate human-like mid-air trajectories that are in accordance with well-established phenomena such as Fitts’ Law (Fitts54) and the Two-Thirds Power Law (Lacquaniti83). RL-based simulation may also provide valuable information for predicting usability- and ergonomics-related criteria and to aid in interface design. For instance, Cheema et al. trained a torque-actuated biomechanical arm model in a mid-air pointing task, and used the model to predict fatigue of real human subjects performing the task (Cheema20). In this paper, we build on these developments and use RL to model interaction in VR.
2.3. Creating Simulated Users
By combining perception models, musculoskeletal models, and physically simulated input devices, we can train RL agents, or simulated users, to model and simulate intricate interaction tasks, such as those requiring visuomotor control. An early example of such control was presented in (Nakada18), where Nakada et al. introduced a virtual human model and used deep learning to learn reaching and tracking tasks.
The development of computationally efficient simulation software, such as MuJoCo (Todorov12), has permitted the use of deep RL for flexible problem formalisations of interaction tasks. This allows a researcher to guide an agent’s learning through reward functions, as was demonstrated in (Ikkala22) by Ikkala et al. with their User-in-the-Box (UitB) approach, where they trained a muscle-actuated model of the upper torso and right arm to complete a series of interaction tasks. UitB allows combining biomechanical models with perception models and movement-based interaction tasks through a modular software structure, which facilitates integrating custom model components. In our work, we exploit this modularity and extend UitB to VR environments.
MyoSuite (caggiano2022myosuite), another user simulator implemented in MuJoCo, focuses on dexterous hand, lower- and full-body movements. MyoSuite also provides a conversion tool for converting OpenSim models into MuJoCo format (MyoConverter [^3] https://github.com/MyoHub/myoconverter (wang2022myosim; Ikkala20)), which can also be readily added into the UitB framework.
Simulated users could promote ability-based design in VR, as diverse user populations, i.e., populations with different physical or cognitive properties, could be simulated without the risk of physical harm (murraysmith2022simulation; QueBolWil2011; quek2013phd). However, existing approaches that create and evaluate simulated users in the context of HCI are limited to “mock-up” interaction environments that require a re-implementation of the application dynamics real users face during interaction (Fischer21; Ikkala22; Hetzel21). These approaches are thus fundamentally unsuitable for automated biomechanical testing of a given VR prototype. To the best of our knowledge, our work is the first to create simulated users in an ecologically valid VR context by training biomechanical user models directly in the same VR environment that human users interact with.
3. VR Simulation Alignment
As depicted in Figure 1 (right), movement-based interaction in VR can be regarded as a continuous-time closed loop between the two core components of interaction: the user and the VR application. During this interaction, the user obtains multi-sensory information from their environment and produces a neuromuscular control signal to generate body movements, hence manipulating the input device to complete the given interaction task. The VR application, on the other hand, conveys an interactive, virtual environment that can be manipulated via one or multiple input devices. The state of the input device is actively sensed by the VR system, measuring, e.g., how the VR controller is moved, or whether a physical controller button is pressed. As a final step to close the loop, the VR application provides sensory output to the user, such as a rendered image displayed on an HMD, or haptic output in the form of temporary controller vibrations.
This interaction loop not only characterises the movement-based interaction between a real user and a given VR system, but can also be used as a framework for creating model-based simulations of VR interaction. To ensure that the simulated interaction loop acts as a reasonable proxy for real user interaction, i.e., to ensure external validity of the predictions made, the two interaction loops need to be aligned.
As described in detail below, we categorise VR simulation alignment along six dimensions, three of which focus on the alignment of the simulated user and three on the alignment of the simulated VR system.
Physical Alignment
The physics of the interaction, i.e., the biomechanics of the user’s body and the dynamics of the input and output devices, need to be captured adequately by the simulation.
Input Channel Alignment
The simulation needs to sense the state of the input device (e.g., the location and orientation of the VR controller) in the same way as the real VR system. In particular, it has to account for coordinate frame transformations and update rates (i.e., the measuring frequency).
Application Alignment
The simulated VR application should have the same internal states as the target VR application (e.g., it needs to provide the same virtual objects and affordances) and show the same responsive behaviour, given that both applications are provided the same input signals.
Output Channel Alignment
The simulated VR system needs to provide the same output signals (e.g., HMD image) as the real VR system, given that both are in the same state. This ensures that the simulated user has the chance to perceive the same information from the application as a real user.
Perceptual Alignment
The simulated user should perceive their environment (including their own body, physical input devices and the VR application) in the same way as a real user would. This requires appropriate modelling of all sensory modalities relevant for the considered interaction, including vision, proprioception and haptics.
Cognitive Alignment
Given an interaction task (e.g., to successfully complete a specific game level as fast as possible), the simulated user should have a similar internal representation and select the same control output when given the same sensory input. In the context of reinforcement learning, this requires defining an appropriate task-specific reward function such that an optimal control policy resembles real user behaviour.
As the behaviour of the user and the VR application depend on each other in a closed loop manner, errors can easily cascade in time. This results in an increasing divergence between the simulation and the real VR interaction, once an error is introduced. For example, in the Whac-A-Mole game from Section 5, a different hammer grip (e.g., due to a slight offset in the game dynamics) may result in fewer/more/different targets being hit, which has an impact on the movements generated by the simulated user at later times, leading to even greater differences in target hits, and so on. The same holds for a missing or incorrectly implemented game constraint: if the simulated user has to hit a target with less power or less often to receive the same game reward, they will learn a totally different control policy, leading to even more divergent application states. Depending on the complexity and the number of discontinuities in the interaction dynamics, these errors may pollute any variable of interest obtained from the simulation. It is thus essential to ensure alignment in as many dimensions as possible.
4. SIM2VR: Enabling Biomechanical User Simulations in VR
In this section, we propose sim2vr, a system that allows two simulators, the simulated user and the original VR application, to run in parallel while being synchronised temporally and spatially. The aim of this system is to establish a closed interaction loop between the simulated user and the VR application (see Figure 2) that closely resembles the interaction loop between the real user and the VR application. While aligning the behaviour of the simulated user with that of real users in all dimensions constitutes a highly complex task, sim2vr addresses the system-related alignment problems (input channel, application, and output channel alignment) described in Section 3.
Figure 2. To obtain valid and reliable predictions of interactive user behaviour in VR, a simulated user needs to perceive and control exactly the same VR application as real users. Our sim2vr system addresses this by establishing a closed loop between the simulation process and the desired VR application, ensuring that the same input and output signals are generated as with real hardware. The resulting time-continuous integration allows to train and evaluate a biomechanical user simulation directly in a given VR environment.
[…]
A closed-loop between the simulated user (shown as biomechanical model with VR controller and HMD) and the VR application (visualised by two example frames from the Whac-A-Mole game) is established through the SIM2VR system (visualised by an arrow from the simulated user to the VR application and an arrow in the opposite direction, with SIM2VR logo in between). This loop consists of four steps (the following texts are annotated in corresponding text boxes): 1) [Simulated User] ”Select muscle control signals to generate body movement and update position and orientation of VR controller and HMD.” 2) [SIM2VR] ”Send simulated sensor input to the VR application.” 3) [VR Application] ”Update game physics and game logic, compute new reward signal and render the scene.” 4) [SIM2VR] ”Send rendered image, game reward and stateful information to the simulated user.”
The core idea is as follows: the simulated user generates muscle control signals based on their sensory perception of the body, the input devices, and the VR application (e.g., visual information of the scene and proprioceptive information of the body posture). These control signals are used to actuate the muscles, driving the movement of the body, and, consequently, the simulated states of the hardware devices. The position and rotation of the VR controllers and HMD are then measured by virtual sensors and sent to the VR application. The VR application processes this input, computes a reward based on the updated application state (e.g., the current game score) and renders the image perceived through the virtual HMD. Finally, the rendered image is provided to the simulated user, along with the reward and optional “stateful” information (such as time remaining in an interaction task), thus closing the interaction loop.
Crucially, this approach ensures application alignment by giving the simulated user access to exactly the same VR application that real users interact with. This is contrary to existing biomechanical frameworks, such as User-in-the-Box (UitB) (Ikkala22) or MyoSuite (caggiano2022myosuite), which are not designed for modelling interaction with real-world VR applications. These frameworks typically run a physics engine that is incapable of capturing the internal states, dynamics, shapes and visuals of a VR application, which fundamentally limits their ability to reliably predict user behaviour. We intentionally refrained from re-implementing the target VR environment in one of these simulation framework for the purpose of this paper, as this would impose considerable implementation and maintenance overheads, and ultimately result in a low-fidelity replica that is inevitably inferior to the VR environment real users are interacting with.
4.1. Implementation Principles & Details
We have implemented sim2vr as an extension of UitB to Unity [^4] https://unity.com/ applications, based on the following principles (outlined in the remainder of this section):
  • Simulated Hardware
  • Standardised VR Interaction Library
  • Direct Access to the VR Environment
  • Synchronised Simulators
We opted to use Unity, as it provides an extensive set of VR/AR features and has a large community of VR developers. The VR application in the simulated interaction loop can be any existing Unity application or a prototype thereof, either built as a standalone app or run within the Unity Editor.
Figure 3. With sim2vr, simulated users have access to exactly the same visual output as humans. The top figure depicts how real users perceive a VR environment (in this case, the Whac-A-Mole game from Section 5), while the bottom figure shows the downsampled RGB-D image that the simulated user perceives.
[…]
Sample renders of the Whac-A-Mole game as shown to real users (top; annotation: ”Real User View (1280x960)”) and the simulated user (bottom; annotation: ”Simulated User View (120x80)”), respectively.
The simulated user can be created within UitB, which requires selecting a biomechanical model implemented in MuJoCo, as well as one or more perception modules. These perception modules define how the simulated user perceives their surroundings, e.g., via visual or proprioceptive signals. To model how users perceive the virtual environment when wearing an HMD, sim2vr adds a novel vision module denoted as UnityHeadset to UitB. This module enables the simulated user to observe the Unity scene through the same Camera game object that renders the scene to real users (see Figure 3). The rendered images are presented to the simulated user as 120$\times$80 pixel RGB-D arrays. The module also allows the visual observation to be filtered and enhanced, for example by omitting certain colour channels or stacking prior observations.
In the UitB framework, the interaction environment and the rewards that incentivise completing the given interaction task need to be defined in a so-called task module. To enable using Unity applications as interaction environments, sim2vr provides a UnityEnv task module on the UitB side, and a SIM2VR Asset on the Unity side. These are connected via ZeroMQ, which creates a communication channel between the two simulators. As UnityEnv handles the communication, it first receives the rendered images from the virtual HMD, before relaying them to the UnityHeadset. Later, after the simulated user’s pose has been updated, UnityEnv sends the HMD and controller position and rotation to the SIM2VR Asset. A detailed overview of the components of sim2vr and their dependencies is given in Figure 4.
Figure 4. sim2vr enables efficient data transfer between the simulated user (implemented in UitB) and the VR application (implemented in Unity). This is achieved by establishing a ZeroMQ connection between the generic UnityEnv task module and the SIM2VR Asset, both provided by sim2vr. The UnityEnv equips the biomechanical model with VR controllers and an HMD. Visual perception of the VR application is received through the UnityHeadset vision module, based on the rendered HMD images sent to the UnityEnv. The SIM2VR Asset also allows application-dependent rewards, such as game scores, to be defined directly in the VR application using the RLEnv Unity script. The rewards can be used by the simulated user to improve their control policy while interacting with the VR application.
[…]
Components of the SIM2VR platform. The ”Simulated User” consists of a ”Biomechanical Model”, a ”Task Module”, a ”Learned Control Policy” and one or multiple ”Perception Modules”, and is implemented in the UitB framework (which makes use of MuJoCo). The ”VR Application” is implemented in Unity. SIM2VR connects these two components through a ZeroMQ connection that allows to transfer sensor input data (symbolised by a VR controller icon) from the ”Simulated User” to the ”VR Application”, and visual perceptions (symbolised by an eye icon) in the opposite direction. This is achieved by adding the ”UnityEnv” task class and the ”UnityHeadset” vision module to the simulated user, and the ”SIM2VR Asset” to the VR application. The latter also provides a template for setting up the application’s ”RLEnv”. Additional arrows are drawn from ”UnityEnv” to ”Biomechanical Model”, from ”UnityEnv” to ”UnityHeadset” and from ”SIM2VR Asset” to ”RLEnv”. The ”RLEnv” block has a dashed border.
The following principles are essential for the functionality of sim2vr:
Simulated Hardware
The simulated user needs to be equipped with VR controllers and an HMD in order to be able to interact with the VR application in the same way as a real user. To this end, the UnityEnv task module augments the biomechanical MuJoCo model of the simulated user with respective hardware objects. [^5] The module currently comes with mesh files of the Meta Quest 1 & 2 (https://www.meta.com/quest/), however, these can be easily replaced with models of other VR equipment. In particular, it ensures that the VR controllers and HMD are rigidly attached to the hands and head of the biomechanical model.
Standardised VR Interaction Library
To achieve input channel alignment, the VR application needs to sense the location and orientation of simulated hardware just as if a physical controller and HMD were used. sim2vr enables such hardware-agnostic input handling by building on the OpenXR [^6] https://www.khronos.org/openxr/. sim2vr requires the Unity application to handle interaction with VR devices using the OpenXR plugin (version 1.5.3 or later). standard.
Direct Access to the VR Environment
By adding the SIM2VR Asset to the desired Unity application, sim2vr gains direct and genuine access to all Unity variables relevant for the interaction. This has several advantages. First, it permits capturing the output signals, such as the rendered image displayed on the HMD, and directly forwards these signals to the simulated user. [^7] While sim2vr is currently limited to the transmission of visual output signals, we plan to support other feedback modalities such as auditory and haptic output. It is thus key in ensuring output channel alignment. Second, it can be used to define application-specific rewards, e.g., based on a game score. This is done in the RLEnv class provided by the SIM2VR Asset, which must be inherited and defined separately for each Unity application to ensure compatibility with RL training. Third, it allows the improvement of the RL training by providing further stateful information (e.g., the time passed in an episode, or the number of completed subtasks) as observations to the simulated user, and to log any Unity variables of interest (e.g., the current performance in terms of specific subtasks).
Synchronised Simulators
The physics engine of the user simulation and the VR application need to be temporally synchronised. sim2vr achieves this by including the current as well as the next timestamp of the user simulation as metadata in the data packages sent to Unity. These timestamps are used to halt the VR application updates until the same amount of time has elapsed as in the user simulation. To avoid computational overhead, the Unity scene is rendered exactly once per UitB time step, and only relevant data is transferred between the two simulators. Faster than real time simulations are achieved by modifying the Unity timescale parameter (in our experiments we ran Unity at 5$\times$ real time). The UnityEnv module can also trigger the reset function of the SIM2VR Asset and vice versa. This is essential to ensure that both the simulated user and the VR application are reset together when necessary, for example, to start a new episode during RL training, or after completing a game level. The simulators are also spatially synchronised by means of aligning the simulators’ coordinate systems and appropriately transforming coordinates between them.
4.2. Creating User Simulations in Unity: A Step-by-Step Guide
In the following, we first describe how sim2vr can be used to generate user simulations for a given Unity application. Then, following the guide, we demonstrate how sim2vr was used to train a simulated user to play a “Beat Saber” style game implemented in the VR Beats Kit, a third party software freely available on the Unity Asset Store. [^8] https://assetstore.unity.com/packages/templates/systems/vr-beats-kit-168243 A more detailed version of this walkthrough is available at the project website https://github.com/fl0fischer/sim2vr.
4.2.1. Importing the Asset
First, the SIM2VR Asset must be imported into the Unity project. After adding the sim2vr prefab as a game object into the desired scene, the SimulatedUser game object needs to be connected to the VR Controllers and Main Camera provided by the OpenXR plugin. The SIM2VR Asset contains Recorder and Logger game objects, which can be used to record game play and log controller and headset poses, as well as further game data, during training and evaluation. Due to the reliance on game-specific data, the Logger must be adapted to each game separately.
4.2.2. Defining the Reward Function
The majority of reward functions that have been used successfully to date have some common features. Typically, they consist of a task-specific component that incentivises reaching the considered goal(s) (e.g., reaching a target position, maintaining a body posture, or moving at a certain speed) and effort costs. Effort costs usually act as a regularisation, i.e., they constrain the set of possible solutions by favouring policies that achieve the goal with less effort. The task-specific reward can be either sparse (i.e., reward is only given after completing the task) or dense (i.e., rewards are continuously given, depending on how “close” the agent is to completing the task). Dense rewards are preferable from a mathematical point of view, as they provide smooth gradients, guiding the RL agent towards the optimal policy. However, defining the “right” dense rewards for a given interaction task, i.e., rewards that incentivise task completion but are not biased towards a particular strategy, can be quite challenging. We provide a reward scaling tool, which can be used as a starting point for defining such rewards; the tool is described in the supplementary material.
In sim2vr, the task-specific reward needs to be calculated and stored within the CalculateReward method of RLEnv. If the VR application under consideration provides a game score, this score can be used directly to define the task-specific reward. The effort costs can be defined when configuring the UitB user model.
4.2.3. Defining the Reset Functionality
The RLEnv class contains methods that need to be implemented to reset the game to an initial state at the end of each episode. This usually includes destroying game objects created during runtime and resetting variables used to compute the reward. All code related to resetting the reward variables should be defined in the InitialiseReward method. Preparations for the next episode, such as choosing a game level, can be defined in the Reset method, or, if these should be called only once when starting the game, in the InitialiseGame method. The end of an episode is indicated by the UpdateIsFinished method.
4.2.4. Defining the Stateful Information (optional)
Since including an application- and task-dependent time feature as stateful information in the observation can help to train the RL agent, the RLEnv class provides a GetTimeFeature method that allows time information to be appended to the observation of the simulated user. Note that in deterministic games, the inclusion of such a time feature may lead to a control policy that exploits this information instead of relying on (visual) observations.
4.2.5. Defining the Simulated User
After preparing the Unity application for running user simulations, an appropriate user model needs to be created in UitB. This mainly involves the choice of a biomechanical user model (including effort cost) as well as perception modules beyond the default UnityHeadset vision module, both of which can be easily selected within the UitB config file. Other optional parameters in the config file include the position and orientation of the HMD relative to a specific body part of the biomechanical user model, arguments to be passed to the Unity application (e.g., setting game level/difficulty), and RL hyperparameters (e.g., network size, time steps, batch size, etc.).
To ensure that all the target positions required by the VR application can be reached by the biomechanical model, we provide a reach envelope tool. This tool allows interactive visualisation of the body-centred positions that a given simulated user can theoretically reach, along with the target positions that occur in a given VR interaction task; the tool is described in the supplementary material.
[…]
(a) Game View
(b) Simulated User
Figure 5. a) In the VR Beats game, the player controls two lightsabers and has to cut through boxes that are moving towards them. b) Still frames from evaluations of trained simulated users playing the VR Beats game, about to hit an incoming box. Top: A simulated user trained with a high effort cost weight ($0.05$) has learned to keep their arms down and rely on wrist movements. Bottom: A simulated user trained with a low effort cost weight ($0.001$) moves their arms considerably more, sometimes resulting in implausible postures.
4.2.6. Example: VR Beats Kit
In VR Beats, players wield lightsabers in both hands and attempts to hit all incoming boxes while dodging walls and avoiding mines (see Figure 5a). Since the biomechanical user models available in UitB cannot move their torso, we trained the simulated user to play only the first 10 seconds of the game, which does not include any walls (or mines). During the first 10 seconds, a total of 4 boxes pass the player, and missing a box causes the episode to end. A successful hit is rewarded with a score between 50-200 points, depending on the speed and direction of the hit. Below we describe the definition of each functionality required by sim2vr.
Reward Function
As a task-specific reward, we used the original (sparse) game score (calculated based on the number of correct and incorrect target hits), as it was easy to implement and not prone to unwanted biases. We also used the default neural effort cost implemented in UitB.
Reset Functionality
For this game, it is sufficient to invoke the existing onGameOverEvent to signal that the episode has ended, destroy the existing game objects, and then invoke the onRestart event to restart the game. The reward is initialised in the Reset method. To check if the episode has terminated, we make use of the existing getIsGameRunning method of the VR_BeatManager instance.
Stateful Information
No additional stateful information was provided to the simulated user.
Simulated User
As VR Beats spawns targets for both hands, we implemented a bimanual UitB user model by converting a bimanual version of the MoBL ARMS model [^9] https://simtk.org/frs/?group_id=657 (Saul14) using the MyoConverter (Ikkala20; wang2022myosim). To speed up the training process, we replaced the muscle actuators with torque actuators attached directly to the joints. As perception modules, we used the default UitB proprioception module and the UnityHeadset vision module provided by sim2vr. The former allows the simulated user to infer joint angles, velocities and accelerations, as well as muscle activations and index finger position. The latter is configured to use the red colour and depth channels of the RGB-D image, and to stack the current observation with a delayed (0.2 seconds prior) visual observation; this configuration allows the control policy to distinguish between left and right arm targets, and to infer target movement.
Results
We trained three simulated users with neural effort cost weights $0.001$, $0.01$ and $0.05$, respectively. These were chosen based on the suggestions provided by the reward scaling tool. These cost weights allow us to explore how the behaviour of the simulated user can vary from an “explorative” or “enthusiastic” player to a more “lazy” player. For instance, the simulated user trained with the highest effort cost weight ($0.05$) learned to do only the bare minimum to move their arms and hit the incoming targets. In particular, they hit the targets by relying mostly on wrist movements, suggesting that the game design could benefit from modifying the target box trajectories to encourage larger arm movements. The simulated users trained with lower effort costs learned to move their arms more, with the lowest effort leading to a policy where the arm movements are exaggerated. Figure 5b shows still frames of simulated users trained with either high or low effort cost.
5. Predicting Interactive User Behaviour with SIM2VR
In this section, we show how sim2vr can be used to predict movement-based interaction in VR and demonstrate its utility in evaluating VR interface design choices. We therefore developed a VR version of the popular, visuomotorically demanding, fast-paced and dynamic arcade game Whac-A-Mole in Unity with different interface variations. This allows us to investigate in a controlled way the extent to which sim2vr can predict the effect of design choices such as difficulty or target area placement. With sim2vr, we train a UitB model to play the game, and then use this model to predict differences in performance, effort and strategies between different game variations. To assess the quality of these predictions, we compare them to reference data from a user study, where real humans were asked to play the same VR game.
(a) Front View with Target Grid
(b) Target Area Placements
Figure 6. (a) In Whac-A-Mole, targets appear randomly at one of nine fixed positions and must be hit with a hammer within one second to score a point. The transparent targets are shown for visualisation purposes only and are not visible during the game. (b) The game allows for three different placements of the target area.
[…]
Top (subcaption: ”Front View of Target Grid”): Front view of the target plane and the target grid in the Whac-A-Mole game, together with the controlled hammer hitting a target. Bottom (subcaption: ”Target Area Placements”): Three target areas (low, mid, high) differing in terms of position and orientation are shown in 3D space together with the controlled hammer.
5.1. Game Design
In the Whac-A-Mole game the goal is to hit targets (the ”moles”) with a hammer to score points. The targets appear randomly on a $3\times 3$ grid as shown in Figure 6a. If a target is successfully hit within one second, it explodes and disappears (see Figure 2) and the game score is increased by one. Otherwise, the target collapses and no points are given. One round of game play lasts for one minute.
We explore three design choices. First, the game has three difficulty levels defining the maximum number of simultaneously displayed targets (easy: 1, medium: 3, hard: 5). The second design choice, target area placement, determines the position and orientation of the target area with respect to the HMD (low, mid, high), as shown in Figure 6b. For the third design choice, we create two versions of the game. In the constrained version, a velocity threshold needs to be exceeded in order to successfully hit a target, whereas in the unconstrained version, no such threshold exists. Further details on the game design are provided in the supplementary material.
5.2. Training Simulated Users
We train a UitB user model to play the different variations of the game. We use MoblArmsWrist, a MuJoCo version of the MoBL ARMS model (Saul14) of the right arm, as the biomechanical model. This model includes shoulder, elbow and wrist joints, and the muscles actuating these joints, resulting in a total of 7 DOFs and 32 muscles. As the model does not actively model neck and head movements, the pose of the virtual HMD is constant throughout the simulation. To account for this limitation, the target area was designed to be of a sufficiently small size to fit within the field of view of the participants, thus minimising the necessity for head movements. Furthermore, we use a non-scaled version of the biomechanical model, as we are not attempting to replicate a specific participant’s movements, but rather aim to demonstrate sim2vr’s utility in generating movements that could be produced by a human participant. Simulating the between-user variability by simultaneously training differently scaled biomechanical models is left for future work.
We use the UnityHeadset vision module provided by sim2vr, along with the pre-implemented UitB proprioception module, to generate observations during training. The UnityHeadset module is configured to include the green colour and depth channels [^10] As the target colour changes from green to red, it is sufficient to use either the green or red channel to infer whether a target is about to disappear., without stacking visual observations. The proprioception module provides observations regarding the biomechanical model, such as the joint positions, velocities, and accelerations, as well as internal muscle activation states.
For training, we use a reward function consisting of multiple components, which are obtained from states of the VR environment, such as game scores, and from states of the biomechanical model, such as muscle fatigue:
(1) $r(s,a)=w_{s}S+w_{c}v_{h}C_{c}+w_{d}C_{d}+w_{e}C_{e}$
The component $S$ is the game score provided by the game; $C_{c}$ is a term to penalise target contacts with too low a hitting velocity in the constrained version (to encourage faster hits); $C_{d}$ penalises the sum of the distances between the hammer and each active target (to incentivise movement toward any target); and $C_{e}$ is an effort term based on the 3CC-r fatigue model (Looft18; Cheema20). The reward components are weighted with $w_{s}=10$, $w_{c}=2.5$, $w_{d}=1$ and $w_{e}=0.1$. The target contact term $C_{c}$ is additionally weighted by the current hammer velocity $v_{h}$.
Simulated users were trained separately for the constrained and unconstrained versions, while keeping the difficulty level fixed to medium and sampling the target area placements randomly at the beginning of each episode. [^11] While it may be reasonable to continue training the user models on either the easy or hard difficulty levels, we observed that the models trained solely on the medium level already generalise well to the remaining difficulties. For each version we train two instances of the simulated user, called SIMu and SIMa. SIMu is trained on the default game dynamics, i.e., the position of the ”mole” is sampled uniformly randomly among the nine possible positions. SIMa, on the other hand, is trained using an adaptive automated curriculum manually implemented in Unity. Such a curriculum can be thought of as a “personal trainer” that creates a customised training plan for the simulated user based on their current strengths and weaknesses.
The learned control policies were evaluated for 1) all three difficulty levels while keeping the target area placement fixed to mid, and 2) for all target area placements while keeping the difficulty level fixed to medium (six configurations in total). Further details on the training and evaluation procedure can be found in the supplementary material. A frame sequence of the simulated user playing the game with the learned control policy is shown in Figure 7.
Figure 7. Still frames from an evaluation of a simulated user playing Whac-A-Mole. The top row shows the simulated user in MuJoCo, while the bottom row is a video capture from the Unity application. In a typical hitting movement, the simulated user gains momentum (left), strikes the target with the virtual hammer (mid) and moves the hammer away from the target area by bending the elbow (right).
[…]
2x3 still frames of a biomechanical user model equipped with the right VR controller and HMD playing the Whac-A-Mole game. The simulated user (top) and the corresponding VR application view (bottom) are shown below each other for three different time frames.
5.3. User Study Method
Participants
In order to compare the simulated user behaviour against real player behaviour, we conducted a user study where we collected game playing data from 18 participants. All but two participants were students and faculty members at Aalto University; one participant was a student at the University of Helsinki, and one participant was not affiliated with either university. All participants were right-handed. The average age was 28.8 years, with a standard deviation of 6.2 years. Eight of the participants identified as female, nine as male, and one as non-binary.
Experimental Design
Twelve participants played the constrained version of the game, while six participants played the unconstrained version. Each participant played the six difficulty $\times$ target area placement configurations that were also used to evaluate the simulated users. For each trial, we recorded the participant’s body and hammer movement. For the three target area placement configurations (with medium difficulty), we also asked participants to report the Borg Rating of Perceived Exertion (Borg RPE) (Borg82) score after each round.
Procedure
The experiments were conducted in an office room. The participants were seated, but not restricted to the chair. Participants were asked to try to avoid bending and rotating their torso (as the simulated users have a fixed torso), and focus on their arm movements. Participants were allowed to train with the medium difficulty level with mid target area placement. This was to reduce learning effects, and to provide the participants a chance to familiarise themselves with the notion of hitting non-physical floating targets. The participants first played the three difficulty configurations in a counterbalanced order, and were then given a rest period of at least 5 minutes, with the option to rest longer if necessary to ensure their right arm was not fatigued. Finally, participants played the three target area placement configurations in a counterbalanced order, and reported the Borg RPE after each round. The experiment length was approximately 30 minutes, and participants were rewarded with a 15 euro gift card to a local restaurant.
5.4. Results
We demonstrate the predictive capabilities of sim2vr by comparing the performance and effort estimates of simulated and real users in the different variations of the game. We also show that sim2vr is capable of predicting specific strategies that users may exploit when game dynamics allow.
Unless otherwise stated, the following results were obtained with the constrained game version.
5.4.1. Performance
Figure 8. The simulated users SIMu and SIMa (left and mid bars) show similar numbers of target hits and misses in the constrained version of the Whac-A-Mole game as the 12 real users from our study on average (right bars). Differences between the three difficulty levels are predicted well.
[…]
Bar plot showing the average number of target hits and misses (y-axis; hits shown as green bars, misses shown as orange bars stacked on top) per difficulty level (x-axis; ”Easy” on the left, ”Medium” in the middle, ”Hard” on the right) and for the two simulated users ”SIM_u” and ”SIM_a” and averaged across all real users (denoted as ”Users”). Figure title: ”Target Hits and Misses”
Figure 9. Both simulated users SIMu and SIMa (highlighted in red) predict target hit rates that are within the between-user variability observed from real players for the constrained version of the Whac-A-Mole game. The simulations also show a similar decrease in hit rate as the game difficulty increases as most real players.
[…]
Scatter plot separately showing the target hit rate (y-axis) per difficulty level (”Easy” shown in blue, ”Medium” shown in orange, ”Hard” shown in green) for the 12 real users and the two simulated users ”SIM_u” and ”SIM_a”, which are highlighted by vertical red stripes (x-axis). The 14 users are sorted in ascending order by the average hit rate across the three levels (08, 04, 06, 09, SIM_u, 01, 07, 11, 12, SIM_a, 10, 03, 02, 05). Figure title: ”Target Hit Rates – Simulated vs. Real Users”
In Whac-A-Mole, the performance (and thus the difficulty) of a game variant can be assessed by the number of target hits and misses users achieve per round. In Figure 8, the mean number of hits and misses from 12 rounds of game play is shown for the two simulated users SIMu and SIMa for each level, along with the means of the 12 real users. For real players (solid, right bars), both target hits and misses increase from the easy to the medium level. However, this trend does not continue when proceeding to the hard level. Instead, the number of hits remains approximately constant (green), while the number of misses (orange) increases. A similar trend can be observed for both SIMu and SIMa (shown as left and mid bars, respectively). Thus, if used early in the design process, the simulated users could have suggested that increasing the task difficulty beyond the medium level would not lead to higher game scores for average users, but only to more target misses.
In Figure 9, the target hit rate is plotted separately for all users in each difficulty level. While there is a considerable variability in the performance of real users, both simulations consistently predict hit rates that lie within this range. In addition, the decrease in hit rate with increasing game difficulty predicted by both simulated users is also observed for almost all real users (the only exceptions are users 9 and 11, who achieved higher hit rates on the hard than on the medium level).
5.4.2. Effort